OpenStudy (anonymous):

A tree house costs $400. You plan to save $10 each week. Your parents will help pay $60 of the cost. Your savings plan can be modeled with the equation S = 10w + 60. Use a table of values to graph the equation in the first quadrant of the coordinate plane. After how many weeks will you be able to afford the tree house? a. 34 b. 35 c. 36 d. 33

OpenStudy (anonymous):

please help Im stuck!

OpenStudy (texaschic101):

well...I don't know about graphing it, but I know this... s = 10w + 60 400 = 10w + 60 400 - 60 = 10w 340 = 10w 340/10 = w 34 = w I know it would take 34 weeks to be able to afford the tree house

OpenStudy (anonymous):

Thank you so so so much! life saver darling! <3